OpenStudy (anonymous):

are the following curves both circles about the origin? 16y^2-400+25x^2=0 and 16x^2-144=9y^2

OpenStudy (unklerhaukus):

circles around the origin have this form \[x^2+y^2=r^2\]

OpenStudy (anonymous):

when i put both of these into a simplified form they are such: 25x^2+16y^2=400 and 16x^2-9x^2=144. these do form the equation you supplied but i wish to know if it valid as both the x and y have values attached to them and are not simply x^2 and y^2

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

I have been told that the first one is a elipses and that the second one is a hyperbola how are these found?

OpenStudy (unklerhaukus):

\[16y^2-400+25x^2=0\]\[16y^2+25x^2=400\]\[\frac{16}{25}y^2+x^2=16\]\[\frac{16}{25}y^2+x^2=4^2\] we see the first equation cannot be made into the form of a circle , because the x and y scale differently in the equation,

OpenStudy (anonymous):

thank you very much.

OpenStudy (unklerhaukus):

 the first equation is centered at the origin but is it an ellipse not a circle

OpenStudy (unklerhaukus):

the second equation\[16x^2-144=9y^2\]can be rearranged to \[16x^2-9y^2=144\] is this the equation of a circle?

OpenStudy (unklerhaukus):

not only do x and y scale differently, but one is negative compared to the other this is a hyperboloid